OpenStudy (anonymous):

Simplify the equation (-6+i)/(-5+i)

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle \frac{i-6}{i-5}}\)

OpenStudy (solomonzelman):

multiply top and bottom times \(i+5\)

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle \frac{(i-6)\color{red}{\times (i+5)}}{(i-5)\color{red}{\times (i+5)}}}\)

Nnesha (nnesha):

don't we have to change the sign of `imaginary term ` ?

OpenStudy (anonymous):

i^2-30/i^2-25?

OpenStudy (solomonzelman):

Nnesha it doesn't matter at all: \(a^2-b^2=(a-b)(a+b)\) will work to cancel the middle term with i to make the denominator a real number in either case.

OpenStudy (solomonzelman):

Jmiliere, only the denominator is correct.

OpenStudy (solomonzelman):

you forgot the MIDDLE TERM in the numerator

OpenStudy (anonymous):

What middle term

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle (i-6)(i+5)=i^2-6i+5i-30=i^2-i-30 }\)

OpenStudy (anonymous):

So it's I-30/i^2-25

OpenStudy (solomonzelman):

And if is \(i\) is here to denote \(\sqrt{-1}\), then you have: \(\large\color{black}{ \displaystyle \frac{(\sqrt{-1}~)^2-i-30 }{(\sqrt{-1}~)^2-25} }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{-1-i-30 }{-1-25} }\)

OpenStudy (solomonzelman):

hope you're good from here

OpenStudy (anonymous):

31+I/24

OpenStudy (solomonzelman):

are you sure that on the bottom is 24 ?

OpenStudy (anonymous):

It would be 26 because of negatives

OpenStudy (solomonzelman):

yes \(\large\color{black}{ \displaystyle \frac{31+i }{26} }\)

OpenStudy (anonymous):

Thank you!

OpenStudy (solomonzelman):

Yw

Nnesha (nnesha):

ahh, I see. Answer would be the same. but i thought (have learned) we should multiply by the conjugate of the denominator which is -i-5.

OpenStudy (solomonzelman):

-i-5 is same as -(i+5)....